# First-return statistics in Henyey–Greenstein scattering: Motzkin polynomials and the Cauchy kernel

### v3

@@ Abstract
In our previous work [Zeller and Cordery 2020 J. Stat. Mech. 2020 063404], we established that first-return probabilities in 1D scattering expand in Catalan and Motzkin numbers. Extending this framework to 3D Henyey–Greenstein scattering requires a Boundary Truncation Factor (BTF): boundary constraints reduce the effective anisotropy from $g^{n}$ to $g^{n}\cdot\mathrm{BTF}(n,g)$. Monte Carlo simulations ($10^{12}$ scattering events) reveal that the BTF follows a Cauchy kernel with width $m_{x}(g)=4g/(1-g)$, amplitude A(g)=1g(1+g)/2A(g)=1-g(1+g)/2, and peak at scattering order $n=2$. This form achieves 1–2% accuracy for $g<2/3$; five parameters capture the entire parameter space ($2\leq n\leq 100$, $0\leq g\leq 0.95$). A first-principles derivation remains open. The closed-form BTF replaces Monte Carlo in iterative inverse problems, enabling orders-of-magnitude speed-up.
@@ 1 Introduction
First-passage problems appear throughout statistical mechanics—in polymer physics, diffusion-limited aggregation, financial mathematics, and queueing theory [1, 2, 3]. A clean example arises in radiative transport: photons entering a scattering medium execute a three-dimensional random walk and may return to the entry boundary. The statistics of this first-passage event encode how stochastic motion interacts with geometric constraint.
Throughout, “first-return” means first passage back to $z=0$ from $z>0$ with negative $z$-velocity—the photon must exit, not graze inward.
@@
where $C(x)=\sum_{n=0}^{\infty}C_{n}x^{n}$ is the Catalan generating function [17] and $C_{n}=(2n)!/(n!(n+1)!)$. This is distribution-free: it depends only on the zigzag structure, not step lengths. Forward-peaked scattering introduces “flat” steps—events preserving the sign of $z$-velocity. The Motzkin extension [14, 15] handles this, giving
@@ 2.2 Motzkin extension for forward scattering
Catalan numbers count Dyck paths—walks restricted to up- and down-steps. Forward scattering introduces flat steps (events preserving the sign of $z$-velocity), requiring the Motzkin extension [14, 15, 16].
@@ 3.3 The Cauchy kernel
Figure 1: The Boundary Truncation Factor versus scattering order $n$ for $g=0.1,0.3,0.5,0.7,0.9$. Points: extracted from Monte Carlo. Curves: Cauchy kernel (equation 13). The width $m_{x}(g)$ increases with $g$; the amplitude $A(g)$ decreases. At $g=0.9$, deviations from Cauchy become visible. (Figures use “path length $m_{s}$” following MC convention; $m_{s}\equiv n$.)
@@ 6.2 Modified Cauchy kernel
The generalization lets the exponent vary:
$K_{\alpha}(x)=\frac{1}{\left[1+x^{2}\right]^{(1+\alpha)/2}}$ (19)
with shape parameter $\alpha$. Setting $\alpha=1$ recovers Cauchy; $\alpha>1$ gives faster tail decay. This “generalized Cauchy” appears in robust statistics [23, 24]. Note that $\alpha$ is a shape parameter, not a Lévy stability index.
The modified BTF becomes
@@
with $A(g)$ and $m_{x}(g)$ unchanged.
@@ 7.1 Open problems
The integer coefficients (2, 4) in the Cauchy kernel are not rounded values but the closest integers compatible with fitting statistics across 10 independent Monte Carlo runs (Section 3.3). This statistical evidence, combined with the appearance of the Cauchy kernel, suggests that a clean derivation may exist, perhaps through first-passage theory for 3D random walks. We leave this as a challenge for future work.
@@ Appendix A Notation summary
An AI-assisted language model (Claude.ai, Anthropic) was used during manuscript preparation to assist in checking and verifying portions of the authors’ own mathematical derivations and to improve clarity of presentation. The AI tool did not generate original results, proofs, or derivations. All mathematics, analyses, and conclusions were independently derived, reviewed, and validated by the authors, who take full responsibility for the content of the manuscript.

### v4

@@ Abstract
We study first-return statistics for photons undergoing three-dimensional Henyey–Greenstein scattering in a semi-infinite medium. In previous work, we showed that one-dimensional first-passage probabilities expand in Catalan and Motzkin generating functions. Extending to three dimensions requires a Boundary Truncation Factor (BTF) that accounts for the restricted angular phase space imposed by the boundary. Extensive Monte Carlo simulations are used to determine the BTF empirically as a function of scattering order and anisotropy. For moderate anisotropy, the BTF is accurately described by a Cauchy kernel with parameters depending only on the Henyey–Greenstein asymmetry factor. This closed-form expression reproduces Monte Carlo results to 1–2% accuracy over a broad range of scattering orders. At higher anisotropy, systematic deviations from the Cauchy form are observed and can be reduced using a one-parameter generalized kernel. The resulting framework provides a computationally efficient mapping from three-dimensional anisotropic transport to one-dimensional combinatorial first-passage theory.
@@ 1 Introduction
First-passage problems appear throughout statistical mechanics—in polymer physics, diffusion-limited aggregation, financial mathematics, and queueing theory [1, 2, 3]. A clean example arises in radiative transport: photons entering a scattering medium execute a three-dimensional random walk and may return to the entry boundary. We consider a semi-infinite medium occupying the half-space $z>0$, with the entry boundary at $z=0$; photons enter at normal incidence. The statistics of this first-passage event encode how stochastic motion interacts with geometric constraint.
Throughout, “first-return” means first passage back to $z=0$ from $z>0$ traveling in the z-z direction—the photon must exit, not graze inward.
@@
where $C(x)=\sum_{n=0}^{\infty}C_{n}x^{n}$ is the Catalan generating function [17] and $C_{n}=(2n)!/(n!(n+1)!)$. This is distribution-free: it depends only on the zigzag structure, not step lengths. Forward-peaked scattering introduces “flat” steps—events with no $z$-direction change. The Motzkin extension [14, 15] handles this, giving
@@ 2.2 Motzkin extension for forward scattering
Catalan numbers count Dyck paths—walks restricted to up- and down-steps. Forward scattering introduces flat steps (events with no $z$-direction change), requiring the Motzkin extension [14, 15, 16].
@@ 3.3 The Cauchy kernel
Figure 1: The Boundary Truncation Factor versus scattering order $n$ for $g=0.1,0.3,0.5,0.7,0.9$. Points: extracted from Monte Carlo. Curves: Cauchy kernel (equation 13). The width $m_{x}(g)$ increases with $g$; the amplitude $A(g)$ decreases. At $g=0.9$, deviations from Cauchy become visible.
@@ 6.2 Modified Cauchy kernel
Setting $\alpha=1$ in the Cauchy form and letting the exponent vary gives the modified BTF:
@@
with $A(g)$ and $m_{x}(g)$ unchanged. Setting $\alpha=1$ recovers the standard BTF; $\alpha>1$ gives faster tail decay. This “generalized Cauchy” form appears in robust statistics [23, 24]. Note that $\alpha$ is a shape parameter, not a Lévy stability index.
@@ 7.1 Open problems
The integer coefficients (2, 4) are not rounded approximations: fitted values across 10 independent Monte Carlo runs are consistent with exactly 2 and 4 within statistical uncertainty. This, combined with the Cauchy kernel form, suggests a clean derivation may exist. We leave this as a challenge for future work.
@@ Appendix A Notation summary
An AI-assisted language model (Claude Opus 4.5, Anthropic, 2025) was used during manuscript preparation to assist in checking and verifying portions of the authors’ own mathematical derivations and to improve clarity of presentation. The AI tool did not generate original results, proofs, or derivations. All mathematics, analyses, and conclusions were independently derived, reviewed, and validated by the authors, who take full responsibility for the content of the manuscript.